Falconmaster:

Three-fourths of a number increased by one-half is equal to fourteen. Find the number. 10 3/8 18 10 1/8 19 1/3

MARC:

\(\frac{3}{4}x+\frac{1}{2}=14\)

MARC:

subtract both sides by 1/2

MARC:

\(\frac{3}{4}x+\frac{1}{2}-\color{red}{\frac{1}{2}}=14-\color{red}{\frac{1}{2}}\)

MARC:

\(\frac{3}{4}x=\frac{27}{2}\)

MARC:

multiply both sides by 4/3

MARC:

\(\color{red}{\frac{4}{3}}\times\frac{3}{4}x=\frac{27}{2}\times\color{red}{\frac{4}{3}}\)

MARC:

\(x=?\) solve it @Falconmaster